Are the following lines parallel, perpendicular, or neither?
L1 with equation x – 4y = 12
L2 with equation 4x + y = 4
Answer #1
Are the following lines parallel, perpendicular, or neither? L1 with equation x – 4y = 12 L2 with equation 4x + y = 4 ---------------- Write in slope-intercept form: L1: y = (1/4)x-3; slope is 1/4 L2: y = -4x+4; slope is -4 ------------- Since (14)*(-4)= 1- the lines are perpendicular. ============== Cheers, Stan H.
answered by: nimrod
Answer #2
First let's find the slope of x – 4y = 12 *[invoke converting_linear_equations "standard_to_slope-intercept", 1, -4, 12, 2, 1] Now let's find the slope of 4x + y = 4 *[invoke converting_linear_equations "standard_to_slope-intercept", 4, 1, 4, 2, 1] Now multiply the slopes of {{{1/4}}} and {{{-4}}} to get {{{(1/4)(-4)=-4/4=-1}}} Since the product of the two slopes is -1, the two slopes are perpendicular
answered by: Tatianna